OpenStudy (anonymous):

rewrite log3^81=4 in exponential form

OpenStudy (mertsj):

What is the base?

OpenStudy (anonymous):

here are the choices 3^4=81 4^3=81 81^1/2=4

OpenStudy (mertsj):

I want to help you understand. What is the base in your problem?

OpenStudy (anonymous):

I dont know?

OpenStudy (mertsj):

Then number written lower than everything else is the base.

OpenStudy (mertsj):

Here is your problem: log, base 3, of 81 is 4

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Now here is something you MUST learn: The log is the exponent

OpenStudy (mertsj):

So if you leave out the stuff in between, your problem say: log is 4

OpenStudy (mertsj):

Or in other words, the exponent is 4

OpenStudy (mertsj):

So now we know that the base is 3 and the exponent is 4 so we write: 3^4

OpenStudy (mertsj):

And we know that 3 raised to the 4th power equals 81 so we write: 3^4=81 and that is the equation in its exponential form.

OpenStudy (anonymous):

Thank you

OpenStudy (mertsj):

yw